OpenStudy (anonymous):

When adding fractions, and you have to find the lowest common denominator. How do you know when to change the numerator? I am really confused!!!!

OpenStudy (radar):

What ever you do to the denominator you do to the numerator otherwise you change the value of the fraction. If you multiply the denominator by 2 then multiply the numerator.\[1/3 + 3/4=\] What is the smallest number that both 3 and 4 will divide into with out a remainder (LCM). In this the LCM is 12 now examine the first fraction, its denominator is 3. What must 3 be multiplied with to equal 12? Your answer is 4. So if I multiply the denominator by 4 I MUST multiply the numerator by 4. The first fraction becomes 4/12. Treat next one in a similar manner.

OpenStudy (anonymous):

Okay, so the next fraction will be 3/12!!!!!

OpenStudy (idily101):

yes

OpenStudy (anonymous):

Okay, so what if i had a question like 1/4 + 3/10 + 2/5.The LCD is 20 right

OpenStudy (idily101):

yup

OpenStudy (idily101):

1/4 would be 1/5 3/10 would be 6/20 and 2/5 would be 8/20

OpenStudy (idily101):

do ypu understand how I got those fractions

OpenStudy (anonymous):

How would 3/10 be 6/20 and 3x6 is 18??? I dont understand!!!

OpenStudy (radar):

No the next fraction in y example was 3/4 we had already decided that 12 would be the LCD for those two fractions. You then multiplied the denominator by 3 and you got your 12.......But you failed to multiply the numerator by 3 !!!!! Remember what you do to the denominator you MUST do to the numerator. Multiplying it by 3 gives you 9/12 not what you got of 3/12 3/12 reduces to 1/4 not 3/4. Look at 9/12 it will reduce to 3/4, see you did not change the value, only the appearance!

OpenStudy (radar):

Khiyah, take a breath and lets do your example. 1/4+ 3/10+2/5. Your choice of 20 was very good, it is the LCM. Convert your fractions with a denominator of 20 WITHOUT CHANGING THEIR VALUE. 1/4: To get a value of 20 I must multiply denominator by 5, what must I multiply the numerator by 5. 1/4=5/20 3/10 a 2 is used to get 20 use the 2 for the numertator also. 3/10 = 6/20. Do you get the procedure then you should get 8/20 for the last fraction. Please note we did not change the value of the fraction.